OpenStudy (anonymous):

What is the derivative of f(x)=g{f(2x)}

OpenStudy (anonymous):

I know I need to use the chain rule here, I'm just not good at it

OpenStudy (freckles):

\[\text{ hint } \\ \frac{d}{dx}h(f(g(x)))=g'(x) \cdot f'(g(x)) \cdot h'(f(g(x))\]

OpenStudy (freckles):

I have exactly as many functions as you do in yours

OpenStudy (freckles):

so that should be easy to follow

OpenStudy (freckles):

start with your most inner function there and follow what I have

OpenStudy (anonymous):

f'(x)=g'(f(2x))*f'(2x)*2?

OpenStudy (freckles):

bingo! :) except that first part

OpenStudy (freckles):

oh no sorry

OpenStudy (freckles):

bingo on the whole thing didn't really that thing was called f(x)

OpenStudy (anonymous):

Okay, great! Thank you!